OpenStudy (anonymous):

use the factor theorem and synthetic division to decide whether or not x-4 is a factor o f x^3-2x^2+x-1.

OpenStudy (anonymous):

plug in 4 to see if the function evaluate to zero

OpenStudy (anonymous):

synthetic division. write the coefficients 1 -2 1 -1

OpenStudy (anonymous):

put a 4 on the side because you are dividing by x - 4 1 -2 1 -1 4 ________________________________________________

OpenStudy (anonymous):

bring down the 1 1 -2 1 -1 4 _________________________________________________ 1

OpenStudy (anonymous):

4 times 1 is 4 1 -2 1 -1 4 4 _________________________________________________ 1

OpenStudy (anonymous):

-2+4=2 1 -2 1 -1 4 4 _________________________________________________ 1 2

OpenStudy (anonymous):

4*2=8 1 -2 1 -1 4 4 8 _________________________________________________ 1 2

OpenStudy (anonymous):

8+1=9 1 -2 1 -1 4 4 8 _________________________________________________ 1 2 9

OpenStudy (anonymous):

4*9=36 1 -2 1 -1 4 4 8 36 _________________________________________________ 1 2 9

OpenStudy (anonymous):

-1+36=36 1 -2 1 -1 4 4 8 36 _________________________________________________ 1 2 9 35

OpenStudy (anonymous):

so the answer is no, it does not divide evenly. it leaves a remainder of 35

OpenStudy (anonymous):

and in fact you now know that \[\frac{x^3-2x^2+x-1}{x-4}=x^2+2x+9+\frac{35}{x-4}\]

OpenStudy (anonymous):

or to put it another way, \[x^3-2x^2+x-1=(x-4)(x^2-2x+9) +35

OpenStudy (anonymous):

\[x^3-2x^2+x-1=(x-4)(x^2-2x+9) +35\]